40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 24 15:24:13 2021, gautam, Update, LSC, Notes on tests 
    Reply  Wed Mar 24 22:54:49 2021, gautam, Update, LSC, New day, new problems PRMI3f_noArmssensMat.pdfREFL55_whtGainStepping.pngREFL55_whtGainStepping2.png
       Reply  Mon Mar 29 19:32:46 2021, gautam, Update, LSC, REFL55 whitening checkout REFL55wht.png
Message ID: 15960     Entry time: Wed Mar 24 22:54:49 2021     In reply to: 15958     Reply to this: 15977
Author: gautam 
Type: Update 
Category: LSC 
Subject: New day, new problems 

I thought I'd get started on some of the tests tonight. But I found that this problem had resurfaced. I don't know what's so special about the REFL55 photodiode - as far as I can tell, other photodiodes at the REFL port are running with comparable light incident on it, similar flat whitening gain, etc etc. The whitening electronics are known to be horrible because they use the quad LT1125 - but why is only this channel problematic? To describe the problem in detail:

  • I had checked the entire chain by putting an AM field on the REFL 55 photodiode, and corroborating the pk-to-pk (counts) value measured in CDS with the "nominal" setting of +18dB flat whitening gain against the voltage measured by a "reference" PD, in this case a fiber coupled NF1611.
  • In the above test, I confirmed that the measured signal was consistent with the value reported by the NF1611.
  • So, at least on Friday, the entire chain worked just fine. The PRMI PDH fringes were ~6000cts-pp in this condition.
  • Today, I found that while trying to acquire PRMI lock, the PDH fringes witnessed in REFL55 were saturating the ADC. I lowered the whitening gain to 0 dB (so a factor of 8). Then the PDH fringes were ~20,000cts-pp. So, overall, the gain of the chain seems to have gone up by a factor of ~25. 
  • Given my NF1611 based test, the part of the chain I am most suspicious of is the whitening filter. But I don't have a good mechanism that explains this observation. Can't be as simple as the input impedance of the LT1125 being lowered due to internal saturations, because that would have the opposite effect, we would measure a tiny signal instead of a huge one

I request Koji to look into this, time permitting, tomorrow. In slightly longer term, we cannot run the IFO like this - the frequency of occurrence is much too high and the "fix" seems random to me, why should sweeping the whitening gain fix the problem? There was some suggestion of cutting the PCB trace and putting a resistor to limit the current draw on the preceeding stage, but this PCB is ancient and I believe some traces are buried in internal layers. At the same time, I am guessing it's too much work to completely replace the whitening electronics with the aLIGO style units. Anyone have any bright ideas?


Anyway, I managed to lock the PRMI (ETMs misaligned) using REFL165I/Q. Then, instead of using the BS as the MICH actuator, I used the two ITMs (equal magnitude, opposite sign in the LSC output matrix).

  • The digital demod phase in this config is different from what is used when the arm cavities are in play (under ALS control). Probably the difference is telling us something about the reflectivity of the arm cavity for various sideband fields, from which we can extract some useful info about the arm cavity (length, losses etc). But that's not the focus here - the correct digital demod phase was 11 degrees. See Attachment #1 for the sensing matrix. I've annotated it with some remarks.
  • The signals appear much more orthogonal when actuating on the ITMs. However, I was still only able to null the MICH line sensed in the PRCL sensor to a ratio of 1/5 (while looking at peaks on DTT). I was unable to do better by fine tuning either the digital demod phase, or the relative actuation strength on each ITM
  • The PRCL loop had a UGF of ~120 Hz, MICH loop ~60 Hz.
  • With the PRMI locked in this config, I tried to measure the appropriate loop gain and sign if I were to use the REFL55 photodiode instead of REFL165 - but I didn't have any luck. Unsurprising given the known electronics issues I guess...

I didn't get around to running any of the other tests tonight, will continue tomorrow.


Update Mar 26: Attachments #2 and #3 show that there is clearly something wrong with the whitening electronics associated with REFL55 channels - with the PSL shutter closed (so the only "signal" being digitized should be the electronics noise at the input of the whitening stage), the I and Q channels don't show similar profiles, and moreover, are not consistent (the two screenshots are from two separate sweeps). I don't know what to make of the parts of the sweep that don't show the expected "steps". Until ndscope gets a log-scaled y-axis option, we have to live with the poor visualization of the gain steps which are dB (rather than linearly) spaced. For this particular case, StripTool isn't an option either because the Q channel as a negative offset, and I opted agains futzing with the cabling at 1Y2 to give a small fixed positive voltage instead. I will emphasize that on Friday, this problem was not present, because the gain balance of the I and Q channels was good to within 1dB.

Attachment 1: PRMI3f_noArmssensMat.pdf  140 kB  Uploaded Thu Mar 25 00:15:27 2021  | Hide | Hide all
PRMI3f_noArmssensMat.pdf
Attachment 2: REFL55_whtGainStepping.png  62 kB  Uploaded Fri Mar 26 17:22:41 2021  | Hide | Hide all
REFL55_whtGainStepping.png
Attachment 3: REFL55_whtGainStepping2.png  61 kB  Uploaded Fri Mar 26 17:22:46 2021  | Hide | Hide all
REFL55_whtGainStepping2.png
ELOG V3.1.3-